Im not sure if this is under probabality and statistics but here is my problem.
If there are 2.5 times as many female as there are male. How many males are there if there are 175 people?
Please provide me with a solution. Thanks.

Answer by edjones(8007)

so there is 1 male for every 2.5 females.
2.5+1=3.5
Let m=number of males
m/175=1/3.5 Proportion
m=175/3.5
m=50
.
Ed